OpenStudy (anonymous):

Write an equation for a line perpendicular to 3y-15x=12 and passing through the point (-10,-2) Y=?? plz help

OpenStudy (anonymous):

the slope would be \[1\over15x\]

OpenStudy (whpalmer4):

First step is to find the slope of the line to which you want to make your line perpendicular. You can find that by rearranging the equation into the slope-intercept form: \(y=mx+b\) and then you read the slope out (it is the value of \(m\)). Second step is to find the slope of the line you are creating. The product of the slopes of two perpendicular lines is always -1. So, your slope for your new line will be \(\dfrac{-1}{m}\) where \(m\) is the slope we obtained in step 1. Finally, use the formula for a line given a known point \((x_0,y_0))\) to construct your equation. The formula is \[y - y_0 = m(x-x_0)\] Note that this value of \(m\) is the one we found in step 2, not the first one.